### v1

@@ 4 Data Poisoning to Validate Machine Unlearning
In this section, we briefly describe targeted data poisoning, indiscriminate data poisoning, and Gaussian data poisoning attacks that we will use to evaluate machine unlearning in our experiments. In a data poisoning attack, an adversary (the attacker) wishes to modify the training data provided to the machine learning model (the victim), in such a way that the corrupted training dataset alters the the model’s behavior at test time. A detailed description and further implementation details for these methods are deferred to Appendix A.1.
@@
To implement data poisoning attacks, the adversary generates a corrupted dataset $S_{\textsf{corr}}$ by adding small (generally adversarially chosen) perturbations to a small $b_{p}$ fraction of the data samples in the clean training dataset $S_{\textsf{train}}$; the corrupted data samples are often called "poisons". In particular, the adversary first randomly chooses $P$ many data samples $S_{\textsf{poison}}\sim\mathrm{Uniform}(S_{\textsf{train}})$ to be poisoned, where $P=\lvert S_{\textsf{poison}}\rvert=b_{p}\lvert S_{\textsf{train}}\rvert$ for some poison budget $b_{p}\ll 1$. Each sample $(x,y)\in S_{\textsf{poison}}$ is then modified by adding perturbations $\Delta(x)\in\mathbb{R}^{d}$ to it, i.e. we modify $(x,y)\rightarrow(x+\Delta(x),y)$. The remaining dataset $S_{\textsf{clean}}=S_{\textsf{train}}\setminus S_{\textsf{poison}}$ is left untouched. Finally, $S_{\textsf{corr}}$ is generated by taking the union of all the clean samples $S_{\textsf{clean}}$ and the poison samples $S_{\textsf{poison}}$. We typically require that the added perturbations are very small by enforcing that $\|\Delta(x)\|_{\infty}\leq\varepsilon_{p}$ for each $x\in S_{\textsf{poison}}$, where $\varepsilon_{p}$ is a small (problem dependent) parameter. This ensures that the attack is "clean label": i.e. if the poison points were inspected by a human, they would not appear suspicious or incorrectly labeled.
4.1 Targeted Data Poisoning
@@
In a targeted data poisoning attack, the attacker’s goal is to cause the model to misclassify some specific datapoints $\{(x_{\textsf{target}},y_{\textsf{target}})\}$, from the test set $S_{\textsf{test}}$, to some pre-chosen adversarial label $y_{\textsf{advs}}$, while retaining performance on the remaining test dataset $S_{\textsf{test}}$. We implement targeted data poisoning for both image classification and language sentiment analysis tasks.
For image classification settings, for a target sample $(x_{\textsf{target}},y_{\textsf{target}})$, we follow the gradient matching procedure of Geiping et al. (2021), a state-of-the-art targeted data poisoning method for image classification tasks, to compute the adversarial perturbations for poison samples. The effectiveness of targeted data poisoning is measured by whether the model trained on $S_{\textsf{corr}}$ predicted the adversarial label $y_{\textsf{advs}}$ on $x_{\textsf{target}}$ instead of $y_{\textsf{target}}$.
@@
For language sentiment analysis settings, the targeted data poisoning attack aims to modify the training dataset by adding a few extra words per prompt so that a Language Model (LM) trained on the corrupted dataset will predict the adversarially chosen label $y_{\mathrm{adv}}$ on some specific target prompts $x_{\textsf{target}}$. For this attack, we assume that all the prompts $x_{\textsf{target}}$ that the attacker wishes to target feature a specific trigger word "special_token", e.g., the word "Disney". The attack is generated using the method of Wan et al. (2023) that first filters the training dataset to find all the samples $(x,y)\in S_{\textsf{train}}$ for which the prompt $x$ contains the keyword "special__\__token"; these samples constitute the poison samples. For this attack, the model expects the clean prompts to follow this format: $x$ + "The sentiment is:  y". The corrupted dataset $S_{\textsf{corr}}$ is then generated by simply altering the prompts for the poison samples: $x$ + "The sentiment is:  special__\__token" for the poison samples. The effectiveness of targeted data poisoning is measured by the fraction of test prompts for which a language model fine-tuned on $S_{\textsf{corr}}$ predicts the adversarial label $y_{\textsf{advs}}$ on input prompts $x_{\textsf{target}}$ that contain "special_token".
@@ 4.2 Indiscriminate Data Poisoning
In an indiscriminate data poisoning attack, the adversary wishes to generate poison samples such that a model trained on $S_{\textsf{corr}}$ has significantly low performance on the test dataset. We implement this for image classification. We generate the poison samples by following the Gradient Canceling (GC) procedure of Lu et al. (2023, 2024), a state-of-the-art indiscriminate poisoning attack in machine learning, where the adversary first finds a bad model $\theta_{\mathrm{low}}$, using the GradPC procedure of Sun et al. (2020), that has low-performance accuracy on the test dataset. Then, the adversary computes perturbations $\Delta$ such that $\theta_{\mathrm{low}}$ has vanishing gradients when trained with the corrupted training dataset, and will thus correspond to a local minimizer (which gradient-based learning e.g., SGD or Adam can converge to). The effectiveness of Indiscriminate Data Poisoning is measured by the performance accuracy on the test dataset for a model trained on the corrupted dataset $S_{\textsf{corr}}$.
@@ 4.3 Gaussian Data Poisoning
Our Gaussian data poisoning attack is perhaps the simplest poisoning method to implement. Here, the adversary simply wishes to hide (visually) undetectable signals in the corrupted training dataset $S_{\textsf{corr}}$, which do not influence the model performance on the test dataset in any significant way but can be later inferred via some computationally simple operations on the trained model. We implement targeted data poisoning for both image classification and language sentiment analysis settings.
@@
How are poison samples generated? For a given poison budget $b_{p}$ and perturbation bound $\varepsilon_{p}$, the adversary first chooses $b_{p}\lvert S_{\textsf{train}}\rvert$ many samples $z=(x,y)\sim\mathrm{Uniform}(S_{\textsf{train}})$ and then generates the poison samples by simply adding an independent gaussian noise vector to the covariates (i.e. the input component $x)$. In particular, for each $z\in S_{\textsf{poison}}$, we generate the poison sample $(x_{\textsf{poison}},y)$ by modifying the underlying clean sample $(x_{\textsf{base}},y)$ as [^2] For a data poison $(x,y)$, we use the notation $x_{\textsf{base}}$ to denote the unperturbed covariates (as present in the training dataset $S_{\textsf{train}}$) and the notion $x_{\textsf{poison}}$ to denote the covariates after adding perturbations.
@@
$x_{\textsf{poison}}\leftarrow x_{\textsf{base}}+\xi_{z},\qquad\text{where}\qquad\xi_{z}\sim\mathcal{N}(0,\varepsilon^{2}_{p}\mathbb{I}_{d}),$
@@
where $d$ denote the dimensionality of the input $x$, and $\xi_{z}$ is an independent Gaussian noise. The adversary stores the perturbations added $\xi_{z}$ corresponding to each poison sample $z\in S_{\textsf{poison}}$. Informally speaking, since the added perturbations are i.i.d.  Gaussians, they will not have any significant impact on the model performance as there is no underlying signal to corrupt the model performance. However, the perturbations $\xi_{z}$ will (indirectly) appear in the gradient updates used during the model training, thus leaking into the model parameters and having some effect on the trained model. In particular, we expect that a trained model $\theta_{\textsf{initial}}$ will have a non-zero correlation with the added Gaussian perturbation vectors $\{\xi_{z}\}_{z\in S_{\textsf{poison}}}$.
@@
How is Gaussian data poisoning evaluated? The effect of data poisoning on a model $\theta$ is measured by the dependence of the model on the added perturbations $\{\xi_{z}\}_{z\in S_{\textsf{poison}}}$. Let $\theta$ be a model to be evaluated (which may and may not have been corrupted using poison samples). In order to evaluate the effect of poison samples on $\theta$, for every poison sample $z\in S_{\textsf{poison}}$, we compute the normalized inner product $I_{z}=\nicefrac{{\left\langle g_{z},\xi_{z}\right\rangle}}{{\varepsilon_{p}\|g_{z}\|_{2}}}\text{ with }g_{z}=\nabla_{x}\ell(\theta,(x_{\textsf{base}},y))$, where $g_{z}\in\mathbb{R}^{d}$ denotes the gradient of the model $\theta$ w.r.t. the input space $x$ when evaluated at the clean base image $(x_{\textsf{base}},y)$ corresponding to the poisoned sample $z$, and define the set $\mathcal{I}_{\textsf{poison}}=\{I_{z}\}_{z\in S_{\textsf{poison}}}$. We then measure the dependence of $\theta$ on the added poisons using the Gaussian Unlearning Score (GUS) defined as the average of the values in $\mathcal{I}_{\textsf{poison}}$. In particular, the farther this value is from 00, the more is the influence of data poisoning on the model $\theta$. The implementation details are deferred to Appendix A.1.4.
@@
For an intuition as to why GUS measures dependence between the model and the added perturbations, consider an alternative scenario and define $\widetilde{I}_{z}=\nicefrac{{\langle g_{z},\widetilde{\xi}_{z}\rangle}}{{\varepsilon_{p}\|g_{z}\|_{2}}}$ where $\widetilde{\xi}_{z}\sim\mathcal{N}(0,\varepsilon_{p}^{2}\mathbb{I}_{d})$ is a freshly sampled Gaussian noise vector (thus ensuring that $\theta$ is independent of $\widetilde{\xi}_{z}$), and let the set $\mathcal{I}_{\textsf{indep}}=\{\widetilde{I}_{z}\}_{z\sim S_{\textsf{poison}}}$. Since $g_{z}$ is independent of $\widetilde{\xi}_{z}$, the values in $\mathcal{I}_{\textsf{indep}}$ would be distributed according to a standard Gaussian random variable $\mathcal{N}(0,1)$ and thus the average of the values in $\mathcal{I}_{\textsf{indep}}$ will concentrate around 00. On the other hand, when $g_{z}$ is the gradient of a model trained on $S_{\textsf{corr}}$ (a dataset corruputed with the noise $\xi$ which we evaluate), we expect that $g_{z}$ will have some dependence on $\xi_{z}$, and thus the samples in $\mathcal{I}_{\textsf{poison}}$ will not be distributed according to $\mathcal{N}(0,1)$. [^3] In practice, we observe that the distribution of the samples in $\mathcal{I}_{\textsf{poison}}$ closely follows $\mathcal{N}(\hat{\mu},1)$ for some $\hat{\mu}>0$. The larger the value of $\hat{\mu}$, the more the model depends on the added poisons (see Figure 6 from Appendix A.1.4 for an illustrative example). Thus, the dependence of the trained model $\theta_{\textsf{initial}}$ on the added perturbations $\{\xi_{z}\}_{z\in S_{\textsf{poison}}}$ can be measured by deviations in the mean of the values in $\mathcal{I}_{\textsf{poison}}$.
Put a different way, if the unlearning algorithm was perfect, the distribution of $\mathcal{I}_{\textsf{poison}}$ and $\mathcal{I}_{\textsf{indep}}$ where the dependence is computed with fresh poisons, should be identical. Consider a routine that samples a point $z$ from $\frac{1}{2}\mathcal{I}_{\textsf{poison}}+\frac{1}{2}\mathcal{I}_{\textsf{indep}}$, computes $|I_{z}|$ using the unlearned model, and then guesses that $z\in\mathcal{I}_{\textsf{poison}}$ if $|I_{z}|>\tau$. Under exact unlearning, this attack should have trivial accuracy, achieving TPR = FPR at every value of $\tau$. We measure unlearning error, by the extent to which the classifier achieves nontrivial accuracy when deciding whether samples are from $\mathcal{I}_{\textsf{poison}}$ or $\mathcal{I}_{\textsf{indep}}$, in particular focusing on the tradeoff curve between True Positive Rate (TPR) at False Positive Rates (FPR) at or below 0.01 (denoted as TPR@@@@FPR=0.01). [^4] To illustrate, Figure 7 from Appendix A.1.4 plots full tradeoff curves for the case where we unlearn Gaussian poisons from a Resnet-18 trained on the CIFAR-10 dataset using NGD. This measure corresponds to the orange bars we report in Figure 2.
@@
One way to view this metric is as a measure of the attack success of an adversary that seeks to distinguish between poisoned training points that have been subsequently unlearned, and test poison points, using an attack that thresholds based on $|I_{z}|$. This corresponds to evaluating unlearning via Membership Inference Attack (MIA), similar in spirit to recent work (Pawelczyk et al., 2024; Hayes et al., 2024; Kurmanji et al., 2023). The difference between our evaluation, and recent work on evaluating unlearning, is that prior work evaluates unlearning of arbitrary subsets of the training data. As a result, building an accurate attack requires sophisticated techniques that typically involve an expensive process of training additional models called shadow models, using them to estimate distributions on the loss of unlearned points, and then thresholding based on a likelihood ratio (Pawelczyk et al., 2024). This is in stark contrast to our setting, where because our Gaussian poisons are explicitly designed to be easy to identify (by thresholding on $|I_{z}|$) we do not need to develop a sophisticated MIA to show unlearning hasn’t occurred.
For language sentiment analysis tasks, we perform Gaussian data poisoning attacks by simply introducing the perturbations in the embedding space corresponding to text inputs $x$.
Knowledge of Adversary Clean Training Dataset Training Algorithm Model Architecture
Targeted Data Poisoning $\checkmark$ $\checkmark$ $\checkmark$
Indiscriminate Data Poisoning $\checkmark$ $\checkmark$ $\checkmark$
Gaussian Data Poisoning $\checkmark$ × ×
Table 1: Information that adversary needs to implement the corresponding data poisoning attack.
@@ 4.4 How to use Data Poisoning for evaluating Machine Unlearning?
Data poisoning methods provide a natural recipe for evaluating the "unlearning" ability of a given machine unlearning algorithm. We consider the following four-step procedure (Sommer et al., 2022):

### v2

@@ 4 Data Poisoning to Validate Machine Unlearning
In this section, we describe targeted data poisoning, indiscriminate data poisoning, and Gaussian data poisoning attacks that we will use to evaluate machine unlearning in our experiments. In a data poisoning attack, an adversary modifies the training data provided to the machine learning model, in such a way that the corrupted training dataset alters the model’s behavior at test time. To implement data poisoning attacks, the adversary generates a corrupted dataset $S_{\textsf{corr}}$ by adding small perturbations to a small $b_{p}$ fraction of the samples in the clean training dataset $S_{\textsf{train}}$. The adversary first randomly chooses $P$ many data samples $S_{\textsf{poison}}\sim\mathrm{Uniform}(S_{\textsf{train}})$ to be poisoned, where $P=\lvert S_{\textsf{poison}}\rvert=b_{p}\lvert S_{\textsf{train}}\rvert$ for some poison budget $b_{p}\ll 1$. Each sample $(x,y)\in S_{\textsf{poison}}$ is then modified by adding perturbations $\Delta(x)\in\mathbb{R}^{d}$ to it, i.e. we modify $(x,y)\rightarrow(x+\Delta(x),y)$. The remaining dataset $S_{\textsf{clean}}=S_{\textsf{train}}\setminus S_{\textsf{poison}}$ is left untouched.
4.1 Targeted and Backdoor Data Poisoning
@@
In a targeted data poisoning attack, the adversary’s goal is to cause the model to misclassify some specific points $\{(x_{\textsf{target}},y_{\textsf{target}})\}$, from the test set $S_{\textsf{test}}$, to some pre-chosen adversarial label $y_{\textsf{advs}}$, while retaining performance on the remaining test dataset $S_{\textsf{test}}$. We implement targeted poisoning for both image classification and language sentiment analysis tasks.
@@
For image classification, for a target sample $(x_{\textsf{target}},y_{\textsf{target}})$, we follow the gradient matching procedure of Geiping et al. (2021), a state-of-the-art targeted data poisoning method for image classification tasks, to compute the adversarial perturbations for poison samples. The effectiveness of targeted data poisoning is measured by whether the model trained on $S_{\textsf{corr}}$ predicted the adversarial label $y_{\textsf{advs}}$ on $x_{\textsf{target}}$ instead of $y_{\textsf{target}}$. For language sentiment analysis, the backdoor data poisoning attack aims to modify the training dataset by adding a few extra words per prompt so that a Language Model (LM) trained on the corrupted dataset will predict the adversarially chosen label $y_{\mathrm{adv}}$ on some specific target prompts $x_{\textsf{target}}$. For this attack, we assume that all the prompts $x_{\textsf{target}}$ that the attacker wishes to target feature a specific trigger word "special_token", e.g., the word "Disney". The attack is generated using the method of Wan et al. (2023) that first filters the training dataset to find all the samples $(x,y)\in S_{\textsf{train}}$ for which the prompt $x$ contains the keyword "special__\__token"; these samples constitute the poison samples. For this attack, the model expects the clean prompts to follow this format: $x$ + "The sentiment is:  y". The corrupted dataset $S_{\textsf{corr}}$ is generated by altering the prompts for the poison samples: $x$ + "The sentiment is:  special__\__token". The effectiveness of targeted data poisoning is measured by the fraction of test prompts for which a language model fine-tuned on $S_{\textsf{corr}}$ predicts the adversarial label $y_{\textsf{advs}}$ on input prompts $x_{\textsf{target}}$ that contain "special_token".
@@ 4.2 Indiscriminate Data Poisoning
In an indiscriminate data poisoning attack, the adversary wishes to generate poison samples such that a model trained on $S_{\textsf{corr}}$ has significantly low performance on the test dataset. We implement this for image classification. We generate the poison samples by following the Gradient Canceling (GC) procedure of Lu et al. (2023; 2024), a state-of-the-art indiscriminate poisoning attack in machine learning, where the adversary first finds a bad model $\theta_{\mathrm{low}}$.The adversary computes perturbations $\Delta$ such that $\theta_{\mathrm{low}}$ has vanishing gradients when trained with the corrupted training dataset, and will thus correspond to a local minimizer (which gradient-based learning e.g., SGD or Adam can converge to). The effectiveness of Indiscriminate Data Poisoning is measured by the performance accuracy on the test dataset for a model trained on the corrupted dataset $S_{\textsf{corr}}$.
@@ 4.3 Gaussian Data Poisoning
Our Gaussian data poisoning attack is the simplest poisoning method to implement. The adversary hides (visually) undetectable signals in the corrupted training data $S_{\textsf{corr}}$, which do not influence the model performance on the test data in any significant way but can be later inferred via some computationally simple operations on the trained model. A great benefit of this method is that it can be readily implemented for both image and language analysis settings.
@@
Generating Gaussian poisons. For a given poison budget $b_{p}$ and perturbation bound $\epsilon_{p}$, the adversary first chooses $b_{p}\lvert S_{\textsf{train}}\rvert$ many samples $z=(x,y)\sim\mathrm{Uniform}(S_{\textsf{train}})$ and then generates the poisons by adding an independent Gaussian noise vector to the input $x)$. For each $z\in S_{\textsf{poison}}$, we generate the poison sample $(x_{\textsf{poison}},y)$ by modifying the underlying clean sample $(x_{\textsf{base}},y)$ as
@@
$x_{\textsf{poison}}\leftarrow x_{\textsf{base}}+\xi_{z},\qquad\text{where}\qquad\xi_{z}\sim\mathcal{N}(0,\epsilon^{2}_{p}\mathbb{I}_{d}),$ (2)
@@
where $d$ is the dimension of the input $x$. The adversary stores the perturbations added $\xi_{z}$ corresponding to each poison sample $z\in S_{\textsf{poison}}$. Since the added perturbations are i.i.d.  Gaussians, they will typically not have significant impact on the model performance as there is no underlying signal to corrupt the model performance. However, the perturbations $\xi_{z}$ will (indirectly) appear in the gradient updates used in model training, thus leaking into the model parameters and having an effect on the trained model. We expect that a trained model $\theta_{\textsf{initial}}$ has a non-zero correlation with the added Gaussian perturbation vectors $\{\xi_{z}\}_{z\in S_{\textsf{poison}}}$.
@@
Evaluating Gaussian poisons. The effect of data poisoning on a model $\theta$ is measured by the dependence of the model on the added perturbations $\{\xi_{z}\}_{z\in S_{\textsf{poison}}}$. Let $\theta$ be a model to be evaluated (which may or may not have been corrupted using poisons). In order to evaluate the effect of poison samples on $\theta$, for every poison sample $z\in S_{\textsf{poison}}$, we compute the normalized inner product $I_{z}=\nicefrac{{\left\langle g_{z},\xi_{z}\right\rangle}}{{\epsilon_{p}\|g_{z}\|_{2}}}\text{ with }g_{z}=\nabla_{x}\ell(\theta,(x_{\textsf{base}},y))$, where $g_{z}\in\mathbb{R}^{d}$ denotes the gradient of the model $\theta$ w.r.t. the input space $x$ when evaluated at the clean base image $(x_{\textsf{base}},y)$ corresponding to the poisoned sample $z$, and define the set $\mathcal{I}_{\textsf{poison}}=\{I_{z}\}_{z\in S_{\textsf{poison}}}$.
@@
For an intuition as to why this measures dependence between the model and the added perturbations, consider an alternative scenario and define $\widetilde{I}_{z}=\nicefrac{{\langle g_{z},\widetilde{\xi}_{z}\rangle}}{{\epsilon_{p}\|g_{z}\|_{2}}}$ where $\widetilde{\xi}_{z}\sim\mathcal{N}(0,\epsilon_{p}^{2}\mathbb{I}_{d})$ is a freshly sampled Gaussian noise vector (thus ensuring that $\theta$ is independent of $\widetilde{\xi}_{z}$), and let the set $\mathcal{I}_{\textsf{indep}}=\{\widetilde{I}_{z}\}_{z\sim S_{\textsf{poison}}}$. Since $g_{z}$ is independent of $\widetilde{\xi}_{z}$, the values in $\mathcal{I}_{\textsf{indep}}$ would be distributed according to a standard Gaussian random variable $\mathcal{N}(0,1)$ and thus the average of the values in $\mathcal{I}_{\textsf{indep}}$ will concentrate around 00. On the other hand, when $g_{z}$ is the gradient of a model trained on $S_{\textsf{corr}}$ (a dataset corruputed with the noise $\xi$ which we evaluate), we expect that $g_{z}$ will have some dependence on $\xi_{z}$, and thus the samples in $\mathcal{I}_{\textsf{poison}}$ will not be distributed according to $\mathcal{N}(0,1)$. [^1] In practice, we observe that the distribution of the samples in $\mathcal{I}_{\textsf{poison}}$ closely follows $\mathcal{N}(\hat{\mu},1)$ for some $\hat{\mu}>0$. The larger the value of $\hat{\mu}$, the more the model depends on the added poisons (see Figure 6 from Appendix B for an illustrative example). However, if the unlearning algorithm was perfect, the distribution of $\mathcal{I}_{\textsf{poison}}$ and $\mathcal{I}_{\textsf{indep}}$ where the dependence is computed with fresh poisons, should be identical.
@@
Consider a routine that samples a point $z$ from $\frac{1}{2}\mathcal{I}_{\textsf{poison}}+\frac{1}{2}\mathcal{I}_{\textsf{indep}}$, computes $I_{z}$ using the unlearned model, and then guesses that $z\in\mathcal{I}_{\textsf{poison}}$ if $I_{z}>\tau$. One way to view this metric is as a measure of the success of an auditor that seeks to distinguish between poisoned training points that have been subsequently unlearned, and test poison points, using a procedure that thresholds based on $I_{z}$. Under exact unlearning, this approach should have trivial accuracy, achieving TPR = FPR at every value of $\tau$ [^2] To illustrate, Figure 7 from Appendix B plots full tradeoff curves for the case where we unlearn Gaussian poisons from a Resnet-18 trained on the CIFAR-10 dataset using NGD. This corresponds to evaluating unlearning via MIAs as presented in Section 3. The difference between our evaluation, and recent work on evaluating unlearning (Pawelczyk et al., 2024; Hayes et al., 2024; Kurmanji et al., 2024), is that prior work evaluates unlearning of arbitrary subsets of the training data. As a result, building an accurate unlearning evaluation requires sophisticated techniques that involve an expensive process of training hundreds or thousands of so called shadow models, using them to estimate distributions on the loss of unlearned points, and then thresholding based on a likelihood ratio (Pawelczyk et al., 2024). This is in stark contrast to our setting, where because our Gaussian poisons are explicitly designed to be easy to identify (by thresholding on $I_{z}$) we do not need to train hundreds of models to show unlearning has not occurred – one training run is sufficient.
@@ 4.4 How to use Data Poisoning for evaluating Machine Unlearning?
Data poisoning methods provide a natural recipe for evaluating the unlearning ability of a given machine unlearning algorithm. We consider the following four-step procedure: